-8(x-3) +12x= 15-(2x +3)


answer: -2 (anyone can show me why the answer is -2 and can give a full break down thank you).

Found 2 solutions by solver91311, scott8148:
Answer by solver91311(24713)





Distribute to remove parentheses. Remember has a coefficient of -1.



Collect like terms:



Add to both sides and add to both sides:



Collect like terms:



Multiply both sides by the reciprocal of the coefficient on the variable, namely



John


Answer by scott8148(6628)

distributing ___ -8x + 24 + 12x = 15 - 2x - 3

collecting terms ___ 4x + 24 = -2x + 12

adding 2x and subtracting 24 ___ 6x = -12

dividing by 6 ___ x = -2